OpenStudy (anonymous):

mg(x2-x1)=1/2k(x2^2-x1^2) can you help me rearrange for x1

OpenStudy (anonymous):

If you start by dividing by mg and multiplying by \[ (x_2^2-x_1^2)\] it'll be a good start ^_^

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

\[(x_2^2-x_1^2)(x_2-x_1)=\frac{1}{2kmg}\] \[(x_2^3-x_1^2 \ x_2 -x_2^2 \ x_1-x_1^3)=\frac{1}{2kmg}\] blegh. This i not a happy polynomial :P

OpenStudy (anonymous):

\[(x_2^3-x_1^2 \ x_2 -x_2^2 \ x_1+x_1^3)=\frac{1}{2kmg}\]

OpenStudy (anonymous):

I don't think you can get x1 by itself....

OpenStudy (anonymous):

maybeee.....

OpenStudy (anonymous):

that is what i thought. but this is for ap physics so it has to, i just couldnt get past this step on my own

OpenStudy (anonymous):

maybe this \[(x_2^3-x_1^2 \ x_2 -x_2^2 \ x_1-x_1^3)=\frac{1}{2kmg}\] \[x_2(x_2^2-x_2 \ x_1-x_1^2 \ )=\frac{1}{2kmg}+x_1^3\] quadrtic formula for the guy in the middle? \[x_2=\frac{x_1±\sqrt{x_1^2+4x_1^2}}{2}\] does that do anything or is even allowed?I forget at this hour :/